### v1

# Enhancing Corpus Callosum Segmentation in Fetal MRI via Pathology-Informed Domain Randomization

Abstract
Accurate fetal brain segmentation is crucial for extracting biomarkers and assessing neurodevelopment, especially in conditions such as corpus callosum dysgenesis (CCD), which can induce drastic anatomical changes. However, the rarity of CCD severely limits annotated data, hindering the generalization of deep learning models. To address this, we propose a pathology-informed domain randomization strategy that embeds prior knowledge of CCD manifestations into a synthetic data generation pipeline. By simulating diverse brain alterations from healthy data alone, our approach enables robust segmentation without requiring pathological annotations.
We validate our method on a cohort comprising 248 healthy fetuses, 26 with CCD, and 47 with other brain pathologies, achieving substantial improvements on CCD cases while maintaining performance on both healthy fetuses and those with other pathologies. From the predicted segmentations, we derive clinically relevant biomarkers, such as corpus callosum length (LCC) and volume, and show their utility in distinguishing CCD subtypes. Our pathology-informed augmentation reduces the LCC estimation error from 1.89 mm to 0.80 mm in healthy cases and from 10.9 mm to 0.7 mm in CCD cases. Beyond these quantitative gains, our approach yields segmentations with improved topological consistency relative to available ground truth, enabling more reliable shape-based analyses. Overall, this work demonstrates that incorporating domain-specific anatomical priors into synthetic data pipelines can effectively mitigate data scarcity and enhance analysis of rare but clinically significant malformations.
Keywords:
Corpus callosum Segmentation Domain Randomization Deep Learning Data Augmentation.
1 Introduction
The corpus callosum (CC) is the brain’s largest commissural fiber tract, connecting the left and right cerebral hemispheres and facilitating interhemispheric communication. Malformations of the CC have been associated with a broad spectrum of cognitive, motor, and behavioral outcomes [1]. Among congenital brain abnormalities, corpus callosum dysgenesis (CCD) is one of the most common, affecting approximately 0.3%–0.7% of the general population, and up to 2–3% of individuals with neurodevelopmental disorders [2]. Diagnosing CCD prenatally is challenging due to its anatomical variability and frequent association with other brain anomalies or genetic syndromes [2]. While ultrasound (US) remains the primary screening tool for CCD, fetal MRI can provide more information, particularly in assessing associated abnormalities and cortical malformations [3]. When combined with super-resolution reconstruction, MRI can produce high-quality 3D volumes that reveal malformations often missed by US, providing crucial information for accurate diagnosis and prognosis [3].
In this context, automated segmentation of fetal brain structures on MR images has become increasingly valuable. It enables rapid, reproducible analysis from in utero MRI, allowing extraction of quantitative biomarkers such as shape, length, and volume, which can support early diagnosis, guide clinical decisions, and inform neurodevelopmental outcome predictions [4, 5]. However, few publicly available fetal MRI datasets exist, and those that exist are largely composed of healthy cases. For example, dHCP [6] includes only healthy subjects, while FeTA [7, 8] provides just eight annotated cases with CCD [9]. Such data imbalance towards healthy anatomy limits the generalization of deep learning models, which often fail on pathological cases when trained primarily on healthy data [8].
As a result, pathology simulation has become a major focus in the field as a means to overcome data scarcity and improve model robustness. Existing pathology-informed augmentation strategies can be broadly categorized into patch-based fusion and learning-based generative models. Patch-based methods, such as CarveMix [10] and PEPSI [11], rely on lesion datasets to transplant annotated abnormalities into healthy scans via cut-and-paste operations. In contrast, learning-based approaches model pathology representations directly, using generative frameworks such as intensity- and texture-driven simulators [12], conditional Generative Adversarial Networks (GANs) [13], or diffusion models for anomaly randomization and pseudo-healthy synthesis [14, 15]. While promising, these methods require annotated pathological data, either for lesion sampling or representation learning, making them difficult to apply in rare disease settings.
To reduce reliance on such annotations, some recent efforts have explored domain randomization (DR) as an alternative strategy. Rather than modeling realistic simulations, DR focuses on variability, through controlled, label-driven transformation, such as intensity and contrast randomization, resolution changes, and artifacts simulation, to improve robustness and generalization to unseen domains [16, 17]. Shang et al. [18] expanded the DR framework of SynthSeg by introducing pathology-informed label-space deformations to simulate ventriculomegaly and spina bifida in fetal cases, while using only healthy annotations. While their method improved performance on the targeted conditions, it was limited to these two pathologies and introduced minor trade-offs in healthy-case performance. Critically, no simulation framework to date has targeted CCD, leaving a significant gap in tools for modeling this common but underrepresented malformation.
In this work, we explore whether prior knowledge of CCD-related anatomical alterations, derived from medical literature and expert input, can be embedded into a DR framework to improve segmentation of pathological fetal brains without requiring real pathological annotations during training. We assess whether incorporating domain-specific priors through label-space augmentations enables models trained solely on healthy data to generalize to different CCD types. Our results show that this approach substantially improves segmentation performance on CCD cases while preserving accuracy on healthy subjects and subjects with other pathologies. Moreover, the resulting segmentations yield anatomically consistent outputs that support robust biomarker extraction; notably, CC length derived from our automated segmentations shows clinically meaningful differences between CCD subgroups.
2 Methods
2.1 Datasets
This study includes a total of 370 fetal subjects sourced from both private clinical cohorts and publicly available datasets. An overview of the datasets is provided in Table 1; we refer readers to the original publications for full details on acquisition protocols and annotation procedures. All data were anonymized and approved by the appropriate ethics committees prior to inclusion.111 KISPI: Ethical Committee of the Canton of Zurich, Switzerland (Decision numbers: 2017-00885, 2016-01019, 2017-00167). CHUV: Approved by the ethics committee of the Canton de Vaud (decision number CERVD 2021 00124). dHCP: Approved by the UK Health Research Authority (REC: 14/LO/1169), with written parental consent. For each subject, we used a super-resolution reconstructed T2-weighted MRI, resampled to an isotropic resolution of 0.5×0.5×0.5mm3 and padded or cropped to a fixed size of 256×256×256 voxels. For the KISPI, dHCP, and STA datasets, annotations were harmonized into a consistent 8-class labeling scheme: cerebrospinal fluid (CSF), grey matter (GM), white matter (WM), ventricles (VM), cerebellum (CBM), deep grey matter (SGM), brainstem (BSM), and CC. Details of the harmonization procedure are provided in the Supplementary Section S1.
Training data. Our segmentation models were trained exclusively on healthy subjects from the KISPI and STA datasets (49 cases total). These datasets provided high-quality segmentations, including the CC.
Healthy test data. For evaluation on healthy brains, we used the dHCP dataset (248 subjects). This dataset was used exclusively for testing, as the CC annotations were found to be inconsistent and unsuitable for training, as discussed in Section 3. Moreover, dHCP serves as an out-of-domain evaluation set to assess the generalization capabilities of our models.
CCD test data. To evaluate segmentation performance on CCD cases, we curated a new dataset, the CHUV-CC dataset, consisting of 26 clinical subjects diagnosed with either complete agenesis (CCA; 12 cases, GA 23–35 weeks) or partial agenesis (pCCA; 14 cases, GA 22–32 weeks). In the pCCA cohort, we have 4 cases of hypoplasia with dysplasia (2 with a kinked morphology, 2 with a stripe-like morphology), 9 cases of hypoplasia (4 without dysplasia and 5 with an anterior remnant change) and 1 case of isolated dysplasia. Each case was reconstructed using one of three SR algorithms—NiftyMIC [19], SVRTK [20], or NeSVoR [21]—selected based on reconstruction quality. Manual segmentations of the corpus callosum were produced by an obstetrician, with more than five years of experience in fetal MRI, using ITK-Snap [22], following the ALBERT annotation protocol [23]. Due to the high cost of manual annotation, only the corpus callosum was labeled in these cases, and only by one rater.
Other test data. To further test generalization to unseen pathologies, we evaluated our models on 47 additional pathological cases from the KISPI cohort, referred to as KISPI-Path. These subjects exhibit a wide range of brain abnormalities, including spina bifida, ventriculomegaly, interhemispheric cysts, and heterotopia, but do not include any isolated CCD cases. This dataset allows us to assess whether models trained to simulate CCD-specific alterations maintain robustness on other fetal brain malformations.
Table 1: Dataset properties. $N_{n}$ – number of neurotypical and $N_{p}$ – pathological subjects.
Site Scanner Acquisition Parameters SR Algorithm Resolution ($mm^{3}$) GA (weeks) $N_{n}$ / $N_{p}$ Annotations
KISPI [24]; [7] GE Signa Discovery MR450 / MR750 (1.5T / 3T) SS-FSE TR/TE: 2500–3500 / 120 ms 0.5×0.5×3.5 mm³ MIAL [25] IRTK [26] $0.5^{3}$ $0.5^{3}$ 20–34 20–34 15/23 16/24 FeTA [24]+ CC [27]
CHUV Siemens MAGNETOM Aera (1.5T) HASTE TR/TE: 1200/90 ms 1.13×1.13×3 mm³ SVRTK [20] NeSVoR [21] NiftyMIC [19] $0.8^{3}$ $0.8^{3}$ $0.8^{3}$ 22–35 25 25 0/24 0/1 0/1 Manual CC Manual CC Manual CC
dHCP [28] Philips Achieva (3T) MB-TSE TR/TE: 2265/250 ms 1.1×1.1×2.2 mm³ IRTK [26]; [29] $0.8^{3}$ 21–38 248/0 dHCP [30]
STA [31] Siemens Skyra / Trio (3T) Philips Achieva (1.5T) SS-FSE TR/TE: 1400–2000/100–120 ms 0.9–1.1×0.9–1.1×2.0 mm³ SR1 [32] $1^{3}$ 21–38 18/0 FeTA [24]+ CC [27]
2.2 Pathology-Informed Domain Randomization for CCD
In this work, we introduce a pathology-informed extension of FetalSynthSeg [33], developed to simulate a wide range of CC anomalies and associated neurodevelopmental alterations, summarized in the Figure 1. Our approach closely follows the original FetalSynthSeg pipeline but modifies its input by applying anatomical augmentations directly to the binary label masks used as segmentation seeds for image synthesis. These deformed segmentations serve both as the basis for generating synthetic intensity images and as training targets for supervised learning of fetal brain tissue segmentation under domain randomization. The augmentation framework is informed by clinical descriptions from Manor et al.[34] and refined through expert input from fetal neuroradiologists. It targets key CCD phenotypes (Figure 1, top), including: complete agenesis (simulated by replacing the CC label with the ventricles), partial agenesis—posterior or anterior (achieved by masking out a proportional segment along the anterior–posterior axis), thinning (implemented with binary erosion using a vertical line structuring element), thickening (through dilation with a spherical kernel and intensity extrapolation), and kinked CC morphology (simulated with a smooth local non-ridid sinusoidal deformation inside the CC’s bounding box along the anterior–posterior axis).
To reflect the frequent co-occurrence of CCD with other central nervous system malformations, we also simulate broader anatomical anomalies. Cortical alterations include thickened cortex (dilation of cortical labels constrained to gray–white matter boundaries), thinned cortex (expansion of CSF within cortical boundaries), and smoothed cortex (sequential dilation and erosion to reduce sulcal folding). For posterior fossa abnormalities, we simulate brainstem and cerebellar hypoplasia (joint erosion of deep brain structures while preserving connectivity to adjacent tissue). Ventriculomegaly is modeled using localized non-rigid radial deformation fields that expand the ventricles, with randomized lateralization, producing unilateral or bilateral presentations.
To ensure plausibility, an expert neuroradiologist reviewed the generated segmentations, both individual alterations and their combinations, to confirm broad anatomical constraints (for example, maintaining realistic ventricular dilation limits and proper continuity in simulations of pontocerebellar hypoplasia). However, the expert did not assess or correct finer-grained details (such as subtle changes in adjacent structures like the cingulate gyrus), both because detailed segmentations of these regions were not available and because enforcing such precision would conflict with the goal of domain randomization.
All these transformations are applied just before the intensity synthesis step and are seamlessly integrated into the FetalSynthSeg pipeline, which ultimately generates synthetic images used to train a 3D U-Net segmentation model for the 8-class tissue segmentation. All augmentations are implemented in 3D and randomized over a range of severity to increase anatomical diversity. The full augmentation code, detailed explanation of the implemented augmentations and parameter ranges, as well as code and documentation, are available at the project GitHub: . We also release a Docker image containing the trained model and its weights to facilitate open science efforts and ensure reproducibility and enable broader testing on external datasets  [^2] Segmentation model is available at the FetalSynthSeg DockerHub page https://hub.docker.com/repository/docker/vzalevskyi/fetalsynthseg/general under ccpathsim tag.
Figure 1: Simulation framework for $C$ alterations (top panel, first row) and broader brain anatomy alterations (top panel, second row), with the bottom panel showing the overall training and evaluation pipeline used in this study.
2.3 Experimental Setup
Evaluated Models. We evaluate four segmentation models to study the effect of domain-randomized pathological augmentation. The two baselines were: BReal, trained exclusively on real fetal MRIs and BSynth, trained on synthetic images generated using the unmodified FetalSynthSeg framework [17]. To assess the impact of incorporating simulated pathologies into the training dataset, we introduced two pathology-augmented variants: SimPath50 and SimPath100, trained on the same synthetic data as BSynth, but with respectively 50% and 100% of training samples modified using our anatomical transformations.
Model architecture and optimization. In all our experiments, we use a 3D-UNet implemented with MONAI [35], that begins with 32 channels and doubles at each upsampling stage. Convolutional layers use 3×3×3 kernels and LeakyReLU activations, except for the final layer, which uses softmax. Skip connections link encoder and decoder levels to preserve spatial resolution. Models were trained with the Adam optimizer (learning rate $10^{-3}$) using a hybrid Dice–Cross Entropy loss. Training stability was managed with a ReduceLROnPlateau scheduler (factor 0.1, patience = 10) and early stopping (patience = 100 iterations). All experiments were conducted on Nvidia RTX 6000 GPUs using PyTorch Lightning with a batch size of 1.
Experiments. First, we evaluate model performance on three test sets: dHCP dataset, CHUV-CC dataset, and KISPI-Path. For dHCP, we report the global average Dice score across the 7 FeTA labels, with the CC merged into white matter in both predictions and ground truth, to account for the unreliable CC annotations in this dataset, as discussed later. For CHUV-CC, we report metrics only for the CC, as it is the only annotated structure. For KISPI-Path, we report metrics for each of the 8 labels individually, as well as the global average. Second, we assess the clinical utility of the resulting segmentations by automatically extracting CC length and evaluating its effectiveness in distinguishing between healthy, pCCA, and CCA cases. This downstream analysis serves as a proxy for the model’s suitability in supporting the extraction of clinically relevant biomarkers.
Model training and architecture. To isolate the effect of synthetic anatomical alterations, we trained the aforementioned series of models on exclusively healthy fetal brain data from the KISPI and STA datasets, using a hybrid Dice Cross Entropy loss, for 8-class brain tissue segmentation. Validation and testing were performed exclusively on real MRI images. All experiments use a five-level 3D U-Net implemented in PyTorch via the MONAI framework, with exactly the same architecture and optimization parameters as in [33].
Metrics and Statistical Analysis. We use standard segmentation evaluation metrics (described in details in the Supp. Section S2.1: generalized Dice Score (DSC) [36], 95th percentile Hausdorff Distance (HD95, reported in mm) [37], and Volume Similarity (VS) [38]. For the CC label, topological correctness was additionally evaluated using the Euler Difference [39], defined as $ED=\left|1-(B_{0}-B_{1}+B_{2})\right|$, where $B_{0}$, $B_{1}$, and $B_{2}$ are the Betti numbers representing the number of connected components, holes, and enclosed cavities, respectively. For the CC, ground truth topology corresponds to a single connected structure with no holes or cavities, and we therefore set the ground truth $B_{0}=1$, $B_{1}=0$, and $B_{2}=0$. For clinical biomarker evaluation, we compute the deviation from normative CC length growth curves as $\Delta_{\text{Growth}}=|SC_{i}-RC_{i}|$ (mm), where $SC_{i}$ is the model’s predicted CC length (estimated as the distance between two furthest points of CC segmentation along the posterior-anterior axis) and $RC_{i}$ is normative values based on the clinical reference from [40] for a given subject at their gestational age $i$. Additional biomarker metrics include absolute CC length error ($\Delta_{\text{Length}}=|CCL_{\text{pred}}-CCL_{\text{gt}}|$). For CCA cases, only $\Delta_{\text{Length}}$ is computed among all metrics, since ground truth segmentations are not available for subjects with CCA. Statistical significance is determined using the Shapiro–Wilk test for normality. When normality holds, paired Student’s t-tests are used; otherwise, Wilcoxon signed-rank tests are applied.
3 Results and discussion
3.1 Effect of Pathology-Informed Augmentations
Table 2: Evaluation of the proposed pathology-informed augmentation models on the dHCP and CHUV-$C$ datasets. For dHCP, Dice and HD95 scores are averaged over the 7 FeTA classes, with $C$ being incorporated into the WM, due to $C$ inconsistencies in its ground truth annotations (see Figure 2). Statistically significant difference between the current model and BSynth is indicated by *
CHUV-CC Metrics BReal BSynth SimPath50 SimPath100 dHCP Metrics BReal BSynth SimPath50 SimPath100
DSC (CC) ($\uparrow$) 0.36* $\pm$ 0.21 0.50 $\pm$ 0.15 0.50 $\pm$ 0.14 0.36* $\pm$ 0.18 DSC (global) ($\uparrow$) 0.73* $\pm$ 0.10 0.83 $\pm$ 0.03 0.83 $\pm$ 0.03 0.84* $\pm$ 0.02
HD95 (CC) ($\downarrow$) 3.52* $\pm$ 1.86 2.87 $\pm$ 1.45 2.32* $\pm$ 1.11 3.78* $\pm$ 2.12 HD95 (global) ($\downarrow$) 25.04* $\pm$ 9.21 8.93 $\pm$ 6.55 8.13 $\pm$ 6.19 5.77* $\pm$ 4.34
VS (CC) ($\uparrow$) 0.73 $\pm$ 0.15 0.80 $\pm$ 0.12 0.81* $\pm$ 0.17 0.68 $\pm$ 0.30 VS (global) ($\uparrow$) 0.82* $\pm$ 0.16 0.97 $\pm$ 0.04 0.98* $\pm$ 0.02 0.98* $\pm$ 0.02
ED (CC) ($\downarrow$) 4.00 $\pm$ 3.59 2.93 $\pm$ 2.23 0.93* $\pm$ 1.27 1.50 $\pm$ 1.45 ED (global) ($\downarrow$) 0.56* $\pm$ 1.42 0.16 $\pm$ 0.65 0.17 $\pm$ 0.50 1.16* $\pm$ 1.73
$\Delta_{\text{Length}}$ ($C$) ($\downarrow$) 7.60 $\pm$ 6.25 6.46 $\pm$ 6.61 6.11 $\pm$ 6.15 7.83 $\pm$ 8.63 $\Delta_{\text{Growth}}$(global)($\downarrow$) 1.73 $\pm$ 1.69 1.63 $\pm$ 1.02 1.66 $\pm$ 1.08 3.12* $\pm$ 2.09
nDSC (CC) ($\uparrow$) 0.45* $\pm$ 0.25 0.67 $\pm$ 0.13 0.65 $\pm$ 0.15 0.44* $\pm$ 0.21
Figure 2: Evaluation on dHCP. Left: Boxplot of Euler characteristic difference and CC length growth curve, comparing our models to the original dHCP annotations. Center: CC length growth trends estimated using quadratic fits for the evaluated models. Right: $C$ segmentations and meshes for representative dHCP subjects, showing improved topology and structure with our models.
KISPI-PATH
Metrics BReal BSynth SimPath100 SimPath50
DSC (global) ($\uparrow$) 0.75
$\pm$0.21
0.75
$\pm$0.15
0.75
$\pm$0.14
0.75
$\pm$0.14
HD95 (global) ($\downarrow$) 4.19
$\pm$5.02
3.01
$\pm$2.27
3.28
$\pm$2.71
3.07
$\pm$2.32
VS (global) ($\uparrow$) 0.96
$\pm$0.05
0.96
$\pm$0.04
0.97
$\pm$0.03
0.97
$\pm$0.02
ED (global) ($\downarrow$) 2.23
$\pm$2.89
1.85
$\pm$3.75
2.76
$\pm$2.48
1.31
$\pm$2.31
Table 3: Evaluation results on the KISPI-Path dataset, averaged across all labels.
Figure 3: CC volume distribution across healthy (dHCP) and CCD cases (left) and CC volume growth chart for dHCP cases (right).
Table 2 summarizes the performance of models trained with and without pathology-informed augmentations. The benefits of incorporating simulated CCD alterations are particularly evident in topology-aware metrics (e.g., ED) and anatomical measures such as $\Delta_{\text{Length}}$. The SimPath50 model consistently outperforms the baselines across all metrics, demonstrating the effectiveness of augmenting healthy data with a moderate proportion of CCD-inspired anatomical variations. In contrast, the SimPath100 model, trained exclusively on simulated pathological anatomies, shows a decline in performance, suggesting that excessive deviation from healthy distributions may hinder generalization. Table 3 as well as supplementary materials Tables S1 and S2 further confirm that SimPath models maintain strong performance on non-CCD pathologies, with no observable degradation relative to the baselines.
On the dHCP dataset, models augmented with pathological simulations perform on par with the baselines, indicating that exposure to synthetic brain alteration does not negatively impact segmentation quality on healthy cases. As shown in Figure 2, all models trained on synthetic data using annotations from [8] achieve significantly better topological consistency than the original dHCP annotations (mean ED = 0.16 for BSynth vs ED = 12.04 for dHCP annotations), and exhibit CC length growth trends closer to clinical reference curves ($\Delta_{\text{Growth}}$ = 1.63 mm for BSynth vs 2.68 mm for dHCP annotations).
Across both CHUV-CC and dHCP test sets, the BSynth model consistently outperforms BReal, which we attribute to the fact that all evaluations are conducted OOD in terms of super-resolution method and acquisition site. The smaller performance gap observed on the KISPI-Path dataset (see Table S1) supports prior findings in [41], emphasizing that SR algorithm and acquisition site are some of the key sources of domain shift in fetal brain MRI.
3.2 Toward Clinically Relevant Biometric Assessment
Figure 4: Evaluation on CHUV-$C$. Top left: Deviation of predicted CC lengths from reference values across three cohorts. Top right: CC length as a function of gestational age on CCA subjects. Bottom: Predicted and reference segmentations and 3D $C$ meshes for a pCCA and CCA case.
Beyond voxel-wise accuracy, we evaluate CC length as a proxy for biometric assessment. Figures 3 and  4 present both qualitative and quantitative results on the CHUV-CC dataset of the CC volume and length, respectively, across healthy and CCD subjects. Figure 3 shows how CC volume drastically decreases with the severity of CC dysgenesis. Furthermore, as can be seen in the Figure 4, the reference $\Delta_{\text{Growth}}$ increases with the severity of the pathology, indicating how CC length values differ from the expected at a given GA. In the top row of Figure 4, both SimPath50 and SimPath100 produce CC length estimates that effectively distinguish between healthy, pCCA and CCA cases, with SimPath100 closely matching the separation achieved by the reference annotations. In contrast, the baseline model struggles to consistently differentiate between pCCA and CCA. These findings indicate that incorporating pathology-informed augmentations enhances the model’s ability to capture clinically relevant anatomical features for downstream analysis and the more severe are the pathologies in the testing dataset the more sever augmentations should be applied for an optimal performance.
4 Conclusions
In this study, we explored whether synthetic data generation can improve fetal brain segmentation in presence of pathology. By embedding anatomical priors into a domain randomization framework, we simulated diverse structural alterations associated with CCD, enabling models trained solely on healthy data to generalize effectively to rare pathological cases.
Our results show that pathology-informed augmentations improve segmentation accuracy in CCD cases, particularly in terms of topological correctness, without compromising performance on normal appearing brains and provided support for clinically meaningful measurements: corpus callosum length, for example, reliably differentiates between healthy, partial, and complete agenesis, suggesting potential for automated severity assessment.
{credits}
4.0.1 Acknowledgements
This research was funded by the Swiss National Science Foundation (182602 and 215641) and ERA-NET Neuron MULTI-FACT project (SNSF 31NE30 203977). We acknowledge the Leenaards and Jeantet Foundations as well as CIBM Center for Biomedical Imaging, a Swiss research center of excellence founded and supported by CHUV, UNIL, EPFL, UNIGE and HUG. This research was also supported by grants from NVIDIA through an Academic Grant Program and utilized the provided NVIDIA RTX6000 ADA GPUs. The Developing Human Connectome Project (dHCP) was funded by the European Research Council (ERC) under the European Union’s Seventh Framework Programme (FP7/2007–2013), Grant Agreement No. 319456.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.